Atoy train travels round one circuit of circular circumference 2.4m in 4.8s calculate
A] average speed?
B] average velocity
1
Expert's answer
2020-11-12T09:34:11-0500

a)


"v=\frac{s}{T}=\frac{2.4}{4.8}=0.5\frac{m}{s}"

b) The  average velocity is zero because the displacement is zero.